Rob and Charles have 36 coins that are nickels and dimes. If the value of the coins is $3.00, how many more dimes than nickels do they have ? Please help me.
Answer by ikleyn(52808)

.
Rob and Charles have 36 coins that are nickels and dimes. If the value of the coins is $3.00,
how many more dimes than nickels do they have ? Please help me.
~~~~~~~~~~~~~~~

d dimes and (36-d) nickels.


Write the "money equation"


    10d + 5*(36-d) = 300  cents.


Simplify and find d


    10d + 180 - 5d = 300

    10d - 5d = 300 - 180

       5d    =     120

        d    =     120/5 = 24.


ANSWER.  24 dimes and  36-24 = 12 nickels.

         24-12 = 12 dimes more than nickels.

Solved.